# Quantum anomalous, quantum spin, and quantum valley Hall effects in pentalayer rhombohedral graphene

### v1

@@ Acknowledgements.
$H_{\text{int}}=\frac{1}{2S}\sum_{{\bm{k}}{\bm{k}}^{\prime}{\bm{q}}}\sum_{ZZ^{\prime}}V_{C}({\bm{q}})c^{\dagger}_{Z}({\bm{k}}+{\bm{q}})c^{\dagger}_{Z^{\prime}}({\bm{k}}^{\prime}-{\bm{q}})c_{Z^{\prime}}({\bm{k}}^{\prime})c_{Z}({\bm{k}}),$ (S33)
where $S$ is the are of the system, and $Z$ and $Z^{\prime}$ represent the spin, valley, sublattice and layer indices. Here, we set the positions of both $K$ and $K^{\prime}$ to be the center in the folded moiré Brillouin zone for simplicity. The form of $V_{C}({\bm{q}})$ is written in the main text. We then construct the Hartree and Fock Hamiltonians as
@@
$\qquad\qquad\qquad\qquad\qquad c^{\dagger}_{Z}({\bm{k}}+{\bm{q}})c_{Z}({\bm{k}}),$ (S34)
@@
$\qquad\qquad\qquad\qquad\qquad c^{\dagger}_{Z}({\bm{k}}+{\bm{q}})c_{Z^{\prime}}({\bm{k}}^{\prime}),$ (S35)
@@
$\qquad\qquad\qquad\qquad\qquad c^{\dagger}_{Z}({\bm{k}}+\bm{G})c_{Z}({\bm{k}}),$ (S36)
@@
$\qquad\qquad\qquad\qquad\qquad c^{\dagger}_{Z}({\bm{k}}^{\prime}+\bm{G})c_{Z^{\prime}}({\bm{k}}^{\prime}),$ (S37)
@@
In the numerical calculation, the summation $\sum_{{\bm{k}}{\bm{k}}^{\prime}\bm{G}}$ is confined to the first and second moiré Brillouin zones. When seeking the ground state of the HF Hamiltonian for a given system parameters, we perform the self-consistent calculations with 10--20 randomized initial states.
@@
Figure S2: (a) The same as Fig. 1. The three paths under consideration are illustrated, labeled by subsequent panels. The charge and spin Chern numbers $(C_{c},C_{s})$ for each region are highlighted in red for QAH, green for QSH, and blue for QVH states. The magenta stars indicate the initial points of each path. (b)(c)(d) Charge and spin Hall conductances $\sigma_{xy}^{c}$ and $|\sigma_{xy}^{s}|$. The coercive field in (b)(c) is represented by $\Delta B$.
S4 Hysteresis
@@
We discuss hysteresis scans of the charge and spin Hall conductances $\sigma_{xy}^{c}$ and $\sigma_{xy}^{s}$. We consider three paths in the $(B_{\perp},B_{\parallel})$ space as illustrated in Fig. S2(a). (In the QSH region of the figure, the spin Chern number $C_{s}$ is denoted as ±$\pm 2$ to reflects the degeneracy of the states #3 and #4 in Table 1. Although either state should be favored due to weak spin-orbit interactions in real systems, we simplify our argument by considering only $|\sigma_{xy}^{s}|$ below.)
@@
Figures S2(b)-(d) depicts anticipated behaviors of the Hall conductance. Figures S2(b) and (c) exhibit hysteresis since transitions from the QAH to QSH/QVH states involve valley flips, leading to a first-order transition. The coercive field $\Delta B$ in both figures is expected to be comparable to the experimentally observed one for the $\nu=1$ QAH effect [5] because the valley of only one band needs to be flipped. Conversely, Fig. S2(d) shows no hysteresis behavior since transition from the QSH to QVH states involves only spin flips. The SU(2) spin-rotational symmetry prevents hysteresis behavior in such cases.
@@
Figure S3: (a) HF band structure at $\nu=3$. The three lowest bands carry $C=+1$ or $-1$. (b) Charge densities $\rho({\bm{r}})$ of the three lowest bands in (a). (c) HF band structure at $\nu=4$. One of the four lowest bands carries $C=0$. (d) Charge densities $\rho({\bm{r}})$ of the four lowest bands in (c).
S5 Numerical results at $\nu=3$ and $4$
@@
Let us discuss the HF band structures and charge densities at $\nu=3$ and 4. Figure S3(a) presents the band structure of the lowest energy state at $\nu=3$. The occupied three lowest bands are separated from other conduction bands. Each band carries the Chern number $C=+1$ or $-1$ depending on the valley. Their charge densities $\rho({\bm{r}})$ are shown in Fig. S3(b). They provide similar (but shifted) extended structure.
@@
Figure S3(c) presents the band structure of the lowest energy state at $\nu=4$. The occupied four lowest bands are separated from other conduction bands. Three bands carry $C=+1$ or $-1$ while the other does $C=0$. Their charge densities are shown Fig. S3(d). The $C=0$ band provides the WLC density structure.

### v2

@@ Appendix C Hartree-Fock calculation
$H_{\text{int}}=\frac{1}{2S}\sum_{{\bm{k}}{\bm{k}}^{\prime}{\bm{q}}}\sum_{ZZ^{\prime}}V_{C}({\bm{q}})c^{\dagger}_{Z}({\bm{k}}+{\bm{q}})c^{\dagger}_{Z^{\prime}}({\bm{k}}^{\prime}-{\bm{q}})c_{Z^{\prime}}({\bm{k}}^{\prime})c_{Z}({\bm{k}}),$ (35)
where $S$ is the area of the system, and $Z$ and $Z^{\prime}$ represent the spin, valley, sublattice and layer indices. Here, we set the positions of both $K$ and $K^{\prime}$ to be the center in the folded moiré BZ for simplicity. The form of $V_{C}({\bm{q}})$ is written in the main text. We then construct the Hartree and Fock Hamiltonians as
@@
$\qquad\qquad\qquad\qquad\qquad c^{\dagger}_{Z}({\bm{k}}+{\bm{q}})c_{Z}({\bm{k}}),$ (36)
@@
$\qquad\qquad\qquad\qquad\qquad c^{\dagger}_{Z}({\bm{k}}+{\bm{q}})c_{Z^{\prime}}({\bm{k}}^{\prime}),$ (37)
@@
$\qquad\qquad\qquad\qquad\qquad c^{\dagger}_{Z}({\bm{k}}+\bm{G})c_{Z}({\bm{k}}),$ (38)
@@
$\qquad\qquad\qquad\qquad\qquad c^{\dagger}_{Z}({\bm{k}}^{\prime}+\bm{G})c_{Z^{\prime}}({\bm{k}}^{\prime}),$ (39)
@@
In the numerical calculation, the summation $\sum_{{\bm{k}}{\bm{k}}^{\prime}\bm{G}}$ is confined to the first and second moiré BZs. When seeking the ground state of the HF Hamiltonian for a given system parameters, we perform the self-consistent calculations with 10--20 randomized initial states. Specifically, we plug random numbers into $\langle c^{\dagger}_{Z^{\prime}}({\bm{k}}^{\prime})c_{Z}({\bm{k}})\rangle$ in the HF Hamiltonian and solve them self consistently.
@@
Figure 6: (a) The same as Fig. 1. The three paths under consideration are illustrated, labeled by subsequent panels. The charge and spin Chern numbers $(C_{c},C_{s})$ for each region are highlighted in red for QAH, green for QSH, and blue for QVH states. The magenta stars indicate the initial points of each path. (b)(c)(d) Charge and spin Hall conductances $\sigma_{xy}^{c}$ and $|\sigma_{xy}^{s}|$. The coercive field in (b)(c) is represented by $\Delta B$.
Appendix D Hysteresis
@@
We discuss hysteresis scans of the charge and spin Hall conductances $\sigma_{xy}^{c}$ and $\sigma_{xy}^{s}$. We consider three paths in the $(B_{\perp},B_{\parallel})$ space as illustrated in Fig. 6(a). (In the QSH region of the figure, the spin Chern number $C_{s}$ is denoted as ±$\pm 2$ to reflects the degeneracy of the states #3 and #4 in Table 1. Although either state should be favored due to weak spin-orbit interactions in real systems, we simplify our argument by considering only $|\sigma_{xy}^{s}|$ below. Note that the QSH state we consider carries nonzero valley Chern number as well [82].)
@@
Figures 6(b)-(d) depicts anticipated behaviors of the Hall conductance. Figures 6(b) and (c) exhibit hysteresis since transitions from the QAH to QSH/QVH states involve valley flips, leading to a first-order transition. The coercive field $\Delta B$ in both figures is expected to be comparable to the experimentally observed one for the $\nu=1$ QAH effect [5] because the valley of only one band needs to be flipped. Conversely, Fig. 6(d) shows no hysteresis behavior since transition from the QSH to QVH states involves only spin flips. The SU(2) spin-rotational symmetry prevents hysteresis behavior in such cases.
@@
Figure 7: (a) HF band structure at $\nu=3$. The three lowest bands carry $C=+1$ or $-1$. (b) Charge densities $\rho({\bm{r}})$ of the three lowest bands in (a). (c) HF band structure at $\nu=4$. One of the four lowest bands carries $C=0$. (d) Charge densities $\rho({\bm{r}})$ of the four lowest bands in (c).
Appendix E Numerical results at $\nu=3$ and $4$
@@
Let us discuss the HF band structures and charge densities at $\nu=3$ and 4. Figure 7(a) presents the band structure of the lowest energy state at $\nu=3$. The occupied three lowest bands are separated from other conduction bands. Each band carries the Chern number $C=+1$ or $-1$ depending on the valley. Their charge densities $\rho({\bm{r}})$ are shown in Fig. 7(b). They provide similar (but shifted) extended structure.
@@
Figure 7(c) presents the band structure of the lowest energy state at $\nu=4$. The occupied four lowest bands are separated from other conduction bands. Three bands carry $C=+1$ or $-1$ while the other does $C=0$. Their charge densities are shown in Fig. 7(d).